# Online metric TSP

### v2

@@ 2 Preliminaries
Though we require that algorithms run in polynomial time, we do not focus on their exact running times. Instead, we evaluate performance via competitive analysis, as is standard in the study of online algorithms. Let $\operatorname{\mathtt{OPT}}(X)$ denote the optimal offline cost for an input sequence $X\in M^{n}$. An online algorithm is said to have competitive ratio $C(n)$, for a function $C\colon\mathbb{N}\to\mathbb{R}$, if its cost is at most $C(n)\operatorname{\mathtt{OPT}}(X)$ for all $n\in\mathbb{N}$ and all $X\in M^{n}$.
@@ 3 Our algorithm
Our algorithm is based on the algorithms for online sorting presented in [1] and [2]. Let us first give a high-level explanation of their approach. Denote by $X^{\prime}$ the set of currently received points. They consider an interval $J$ containing $X^{\prime}$, and partition $J$ into $\sqrt{n}$ subintervals. Specifically, they let $J$ be the smallest interval of the form $[-2^{k},2^{k}]$ containing $X^{\prime}$, increasing $k$ as necessary. Note that $\operatorname{\mathtt{OPT}}(X^{\prime})\in O(2^{k})$, so the subintervals have size $O(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$. Similarly, they partition the array $A$ into $2\sqrt{n}$ subarrays, called blocks. The algorithm generally places the input points into the array, such that no block contains points from different intervals. This ensures, that they only pay a small $O(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$ cost between points inside each block, and pay a large $O(\operatorname{\mathtt{OPT}}(X^{\prime}))$ cost only between blocks. The former cost is paid around $O(n)$ times, and the latter around $O(\sqrt{n})$ times, totaling a cost of $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X^{\prime}))$. It is always possible to place the first half of the points in such a way, after which they consider the remaining empty array cells as one contiguous array, and recursively fill this simulated array with the remaining input points. This leads to a total cost of $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X))$ as the cost per recursion falls geometrically. Each time $k$ is increased, we incur some additional cost of changing the intervals, but this only happens when the bound on $\operatorname{\mathtt{OPT}}(X^{\prime})$ doubles, so again it all sums to $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X))$.
@@
Let us now give a high-level explanation of how we will extend this algorithm from the Euclidean line to a general metric space. There are two main problems to solve: the setting of $J=[-2^{k},2^{k}]$ and bound of $\operatorname{\mathtt{OPT}}(X^{\prime})\in O(2^{k})$ no longer make sense. One might naively try letting $J$ be a ball of radius $2^{k}$, but this cannot generally be covered by $\sqrt{n}$ balls of radius $O(2^{k}/\sqrt{n})$. To see this, think of $d$-dimensional Euclidean space, where $2^{d}$ unit cubes are required to cover a cube of sidelength 2. This is related to why there is a term of size $2^{d}$ in the best-known bound for $d$-dimensional Euclidean space [2]. We instead present an online algorithm maintaining a set of at most $\sqrt{n}$ balls of radius $r=\Theta(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$ covering all received input points. Only when $\operatorname{\mathtt{OPT}}(X^{\prime})$ doubles, we allow changing the radius $r$ and resetting the set. Of course, we don’t know the value of $\operatorname{\mathtt{OPT}}(X^{\prime})$, but it turns out that we can successfully use the minimum spanning tree weight as a polynomial-time computable proxy. This proxy will be essential when showing that our covering only needs $\sqrt{n}$ balls.
@@
Let $(M,d)$ be a metric space and $X\subseteq M$ a finite subset of points. If the metric $d$ is clear from context, we denote by $\operatorname{\mathtt{MST}}(X)$ the total weight of a minimum spanning tree of the complete graph on $X$, where edge weights are given by $d$.
Computing exact metric TSP length is NP-hard, as follows by a straightforward reduction from the undirected Hamiltonian cycle problem, one of Karp’s 21 NP-complete problems [4]. Luckily, there is a simple 2-approximation based on minimum spanning trees [6]. We prove that $\operatorname{\mathtt{MST}}(X)$ is a good proxy of $\operatorname{\mathtt{OPT}}(X)$ using this well-known idea.
@@
Recall that $\operatorname{\mathtt{OPT}}(X)$ is the length of some (shortest) walk covering all points in $X$, so in particular it is the total weight of some connected multigraph on $X$. Since $\operatorname{\mathtt{MST}}(X)$ is the minimum total weight of a connected graph on $X$, we have $\operatorname{\mathtt{MST}}(X)\leq\operatorname{\mathtt{OPT}}(X)$. To show the remaining inequality, consider walking along the face created by placing a minimum spanning tree in the plane. This gives a walk of length $2\operatorname{\mathtt{MST}}(X)$ covering $X$. Since $\operatorname{\mathtt{OPT}}(X)$ is the length of a shortest walk covering $X$, we have $\operatorname{\mathtt{OPT}}(X)\leq 2\operatorname{\mathtt{MST}}(X)$. ∎
The following property of our proxy will be useful.
@@
Let $(M,d)$ be a metric space and $Y\subseteq X\subseteq M$ finite subsets of points. Then $\operatorname{\mathtt{MST}}(Y)\leq\operatorname{\mathtt{OPT}}(X)$.
Proof.
By Lemma 5, we have $\operatorname{\mathtt{MST}}(Y)\leq\operatorname{\mathtt{OPT}}(Y)$, and by the triangle inequality, we have $\operatorname{\mathtt{OPT}}(Y)\leq\operatorname{\mathtt{OPT}}(X)$. ∎
As sketched, our algorithm will cover the input points by a dynamic set of balls. Concretely, we will work with a set of centers.
Definition 7.
Let $(M,d)$ be a metric space and $X\subseteq M$ a set of points. A subset $C\subseteq X$ together with a real number $r$ forms a covering $(C,r)$ of $X$ if for every $x\in X$, there exists a $c\in C$ with $d(c,x)\leq r$. If so, we call each point in $C$ a center and $r$ the radius of the covering. A covering $(C,r)$ is minimal if for every $c,c^{\prime}\in C$ with $c\neq c^{\prime}$, we have $d(c,c^{\prime})>r$.
Note that for a fixed radius, this definition of minimal matches the usual definition that no strict subset satisfies the same defining properties. Minimal coverings will be useful to us, because they are small in terms of our proxy, in the sense of the following lemma.
Lemma 8.
Let $(M,d)$ be a metric space, in which $(C,r)$ is a minimal covering of a set $X\subseteq M$. Then $2\operatorname{\mathtt{MST}}(X)\geq\operatorname{\mathtt{MST}}(C)\geq(|C|-1)r$.
Proof.
Combining Lemmas 6 and 5, we get $2\operatorname{\mathtt{MST}}(X)\geq\operatorname{\mathtt{OPT}}(X)\geq\operatorname{\mathtt{MST}}(C)$. Every edge in the complete graph on $C$ has length at least $r$, and a tree on $|C|$ points has $|C|-1$ edges, so $\operatorname{\mathtt{MST}}(C)\geq(|C|-1)r$. ∎
When receiving a new point $x$, we will generally update our minimal covering by the following simple subroutine.
Increase-Minimal-Covering($C,r,x$) 1. If no $c\in C$ has $d(x,c)\leq r$ then add $x$ to $C$.
This subroutine allows us to maintain a minimal covering, for a fixed radius. We state this as the following lemma, which follows directly from the definition.
Lemma 9.
Let $(M,d)$ be a metric space, in which $(C,r)$ is a covering of a set $X\subseteq M$, and let $x\in X$. Then running Increase-Minimal-Covering$(C,r,x)$ modifies $C$ only by insertion, such that $(C,r)$ becomes a minimal covering of $X\cup\{x\}$.
@@
We now present “half” of our algorithm, namely an algorithm handling the first $\lceil n/2\rceil$ input points. First, we give a quick overview. The algorithm will partition the array $A$ into $2\lfloor\sqrt{n}\rfloor$ blocks, and maintain a covering $(C,r)$ with at most $\lfloor\sqrt{n}\rfloor$ centers. If we get too many centers, we will reset the covering with a larger radius. Each block may be assigned to a center, such that every center is assigned at most one block. Initially, every block is unassigned. Generally, if a newly received point lies within the ball of a center, it will be placed inside the block assigned to that center. The exact algorithm is as follows.
Fill-Most-Blocks($n,A,X$) 1. Let $N_{1}=\lfloor\sqrt{n}\rfloor$ and $N_{2}=2N_{1}$. 2. Partition $A$ into $N_{2}$ subarrays, called blocks, of length at least $\lfloor n/N_{2}\rfloor$. 3. Initialize $r=0$ and $C=\emptyset$. 4. For each point $x$ in the stream $X$: (a) Increase-Minimal-Covering$(C,r,x)$. (b) If $|C|>N_{1}$: i. Unassign all blocks. ii. Set $r=4\operatorname{\mathtt{MST}}(X^{\prime})/N_{1}$, where $X^{\prime}$ is the set of known input points. iii. Set $C=\{x\}$. (c) Let $c\in C$ be a center with $d(x,c)\leq r$. (d) If a full block $B$ is assigned to $c$: i. Unassign $B$ from $c$. (e) If no block is assigned to $c$: i. Assign an unassigned non-full block to $c$. (f) Let $B$ be the block assigned to $c$. (g) Place $x$ in the left-most empty cell of $B$.
@@
Claim. After step 4b, $(C,r)$ is a minimal covering of a set containing $x$.
Proof of claim. Clearly, the claim holds if step 4(b)iii just ran, so let us show, that it also holds before and between these resets of the covering. This follows by induction, where the base case is step 3 producing a minimal covering, and the induction step is step 4a with Lemma 9. $\triangleleft$
@@
Consider first the case where the covering was not reset (steps 4(b)i to 4(b)iii) between $x$ and $x^{\prime}$ being placed. Then $B$ was assigned to the same center $c$ with the same radius $r$ both when $x$ and $x^{\prime}$ were placed. By the triangle inequality and Lemma 6, $d(x,x^{\prime})\leq d(x,c)+d(c,x^{\prime})\leq 2r\leq 8\operatorname{\mathtt{OPT}}(X)/N_{1}$. Since at most $\lceil n/2\rceil$ points are placed, the total cost between such pairs of points is at most $(\lceil n/2\rceil-1)8\operatorname{\mathtt{OPT}}(X)/N_{1}\leq 4n\operatorname{\mathtt{OPT}}(X)/N_{1}$.
Consider now the case where the covering was rebuilt between $x$ and $x^{\prime}$ being placed. Then $d(x,x^{\prime})\leq\operatorname{\mathtt{MST}}(X^{\prime})$ where $X^{\prime}$ is the set of points read at any point after placing both $x$ and $x^{\prime}$. This is a worse bound, but this case can only appear once per block per reset of the covering. So the total cost for such pairs of points is at most $N_{2}\operatorname{\mathtt{MST}}(X^{\prime})$ for each reset, where $X^{\prime}$ is the set of points read when resetting. Let $X_{1}$ be the set of points read at the point of a reset, or at the beginning of the algorithm, and $X_{2}$ be the set of points read at the point of the following reset or when the algorithm terminates. Right before the latter reset, $|C|>N_{1}$ and $r=4\operatorname{\mathtt{MST}}(X_{1})/N_{1}$. This is true even if $X_{1}$ is the set at the beginning of the algorithm, as then $X_{1}=\emptyset$ and $r=0$. So Lemma 8 gives us that $2\operatorname{\mathtt{MST}}(X_{2})\geq\operatorname{\mathtt{MST}}(C)\geq(|C|-1)r\geq N_{1}4\operatorname{\mathtt{MST}}(X_{1})/N_{1}=4\operatorname{\mathtt{MST}}(X_{1})$, which simplifies to $\operatorname{\mathtt{MST}}(X_{2})\geq 2\operatorname{\mathtt{MST}}(X_{1})$. The total cost of this type across all rebuilds is thus no more than the geometric series $N_{2}\sum_{i=0}2^{-i}\operatorname{\mathtt{MST}}(X)=2N_{2}\operatorname{\mathtt{MST}}(X)\leq 2N_{2}\operatorname{\mathtt{OPT}}(X)$, where the last inequality uses Lemma 5.
@@ 4.1 Impossibility of optimality for every fixed metric space
By closely following the proof of Theorem 12, replacing their random input by $\mathcal{X}$, we get that the expected cost of any deterministic algorithm on $\mathcal{X}$ is $\Omega(n)$. We provide a full proof in Appendix A, restating this claim as Lemma 14.
@@
By assumption, $\mathcal{A}$ has competitive ratio $O(\sqrt{n})$ in general, and competitive ratio $O(\log n)$ when $x$ is not in the input. It’s easy to see that $\operatorname{\mathtt{OPT}}(\mathcal{X})\in O(n^{4/5})$, so we have $\operatorname{\mathbb{E}}[\mathcal{X}\mid x\in\mathcal{X}]\in O(n^{4/5}\sqrt{n})$ and $\operatorname{\mathbb{E}}[\mathcal{X}\mid x\not\in\mathcal{X}]\in O(n^{4/5}\log n)$. In total, $\operatorname{\mathbb{E}}[\mathcal{X}]\in O(n^{-2/5}n^{4/5}\sqrt{n}+n^{4/5}\log n)\subseteq o(n)$, contradicting $\operatorname{\mathbb{E}}[\mathcal{X}]\in\Omega(n)$. ∎

### v3

@@ 2 Preliminaries
Though our algorithms run in polynomial time, we do not focus on their exact running times. Instead, we evaluate performance via competitive analysis, as is standard in the study of online algorithms. Let $\operatorname{\mathtt{OPT}}(X)$ denote the optimal offline cost for an input sequence $X\in M^{n}$. An online algorithm is said to have competitive ratio $C(n)$, for a function $C\colon\mathbb{N}\to\mathbb{R}$, if its cost is at most $C(n)\operatorname{\mathtt{OPT}}(X)$ for all $n\in\mathbb{N}$ and all $X\in M^{n}$.
@@ 3 Our algorithm
Our algorithm is based on the algorithms for online sorting presented in [1] and [2]. Let us first give a high-level explanation of their approach. Denote by $X^{\prime}$ the set of currently received points. They consider an interval $J$ containing $X^{\prime}$, and partition $J$ into $\sqrt{n}$ subintervals. Specifically, they assume without loss of generality that $x_{1}=0$, and let $J$ be the smallest interval of the form $[-2^{k},2^{k}]$ containing $X^{\prime}$, increasing $k$ as necessary. Note that $\operatorname{\mathtt{OPT}}(X^{\prime})\geq 2^{k-1}$, so the subintervals have size $O(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$. Similarly, they partition the array $A$ into $2\sqrt{n}$ subarrays, called blocks. The algorithm generally places the input points into the array, such that no block contains points from different subintervals. This ensures, that they only pay a small $O(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$ cost between points inside each block, and pay a large $O(\operatorname{\mathtt{OPT}}(X^{\prime}))$ cost only between blocks. The former cost is paid $O(n)$ times, and the latter $O(\sqrt{n})$ times, totaling a cost of $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X^{\prime}))$. It is always possible to place the first half of the points in this manner, after which they consider the remaining empty array cells as one contiguous array, and recursively fill this simulated array with the remaining input points. This leads to a total cost of $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X))$ as the cost per recursion falls geometrically. Each time $k$ is increased, we incur some additional cost of changing the intervals, but this only happens when the bound on $\operatorname{\mathtt{OPT}}(X^{\prime})$ doubles, so again it all sums to $O(\sqrt{n}\operatorname{\mathtt{OPT}}(X))$.
@@
Let us now give a high-level explanation of how we will extend this algorithm from the Euclidean line to a general metric space. There are two main problems to solve: the setting of $J=[-2^{k},2^{k}]$ and bound of $\operatorname{\mathtt{OPT}}(X^{\prime})\geq 2^{k-1}$ no longer make sense. One might naively try letting $J$ be a ball of radius $2^{k}$, but this cannot generally be covered by $\sqrt{n}$ balls of radius $O(2^{k}/\sqrt{n})$. To see this, think of $d$-dimensional Euclidean space, where $2^{d}$ unit cubes are required to cover a cube of sidelength 2. This is related to why there is a term of size $2^{d}$ in the best-known bound for $d$-dimensional Euclidean space [2]. We instead present an online algorithm maintaining a set of at most $\sqrt{n}$ balls of radius $r=\Theta(\operatorname{\mathtt{OPT}}(X^{\prime})/\sqrt{n})$ covering all received input points. Only when $\operatorname{\mathtt{OPT}}(X^{\prime})$ doubles, we allow changing the radius $r$ and resetting the set. Of course, we don’t know the value of $\operatorname{\mathtt{OPT}}(X^{\prime})$, but it turns out that we can successfully use the minimum spanning tree weight as a (polynomial-time computable) proxy. This proxy will be essential when showing that our covering only needs $\sqrt{n}$ balls.
@@
Let $(M,d)$ be a metric space and $X\subseteq M$ a finite subset of points. When the metric $d$ is clear from context, we denote by $\operatorname{\mathtt{MST}}(X)$ the total weight of a minimum spanning tree of the complete graph on $X$, where edge weights are given by $d$.
Computing exact metric TSP length is NP-hard, as follows by a straightforward reduction from the undirected Hamiltonian cycle problem, one of Karp’s 21 NP-complete problems [4]. Luckily, there is a simple 2-approximation based on minimum spanning trees [6]. This well-known argument yields the following lemma, which tells us that $\operatorname{\mathtt{MST}}(X)$ is a good proxy of $\operatorname{\mathtt{OPT}}(X)$. We provide a proof in Section A.1.
@@
As sketched, our algorithm will cover the input points by a dynamic set of balls. Formally, we will work with a set of centers forming a net in the following standard manner.
Definition 6.
Let $(M,d)$ be a metric space and $X\subseteq M$ a set of points. Let $r\geq 0$ be a real number. A subset $C\subseteq X$ is an $r$-net of $X$ if
  • for every $x\in X$, there exists a $c\in C$ with $d(c,x)\leq r$, and
  • for every $c,c^{\prime}\in C$ with $c\neq c^{\prime}$, we have $d(c,c^{\prime})>r$.
If so, we sometimes refer to each point in $C$ as a center, and to $r$ as the radius of the net.
Nets will be useful to us, because they are small in terms of our proxy, in the sense of the following lemma.
Lemma 7.
Let $(M,d)$ be a metric space, in which $C$ is an $r$-net of a set $X\subseteq M$. Then $(|C|-1)r\leq 2\operatorname{\mathtt{MST}}(X)$.
@@
By the triangle inequality, we have $\operatorname{\mathtt{OPT}}(X)\geq\operatorname{\mathtt{OPT}}(C)$. Combining this with Lemma 5, we get $2\operatorname{\mathtt{MST}}(X)\geq\operatorname{\mathtt{OPT}}(X)\geq\operatorname{\mathtt{OPT}}(C)\geq\operatorname{\mathtt{MST}}(C)$. Every edge in the complete graph on $C$ has length at least $r$, and a tree on $|C|$ points has $|C|-1$ edges, so $\operatorname{\mathtt{MST}}(C)\geq(|C|-1)r$. ∎
When receiving a new point $x$, we will generally update our $r$-net by the following simple subroutine.
Increase-Net($C,r,x$) 1. If no $c\in C$ has $d(x,c)\leq r$ then add $x$ to $C$.
This subroutine allows us to maintain an $r$-net, for a fixed radius $r$. We state this as the following lemma, which follows directly from the definition.
@@
Let $(M,d)$ be a metric space, in which $C$ is an $r$-net of a set $X\subseteq M$, and let $x\in X$. Then running Increase-Net$(C,r,x)$ modifies $C$ only by insertion, such that $C$ becomes an $r$-net of $X\cup\{x\}$.
@@
We now present “half” of our algorithm, namely an algorithm handling the first $\lceil n/2\rceil$ input points. First, we give a quick overview. The algorithm will partition the array $A$ into $2\lfloor\sqrt{n}\rfloor$ blocks, and maintain an $r$-net $C$ with at most $\lfloor\sqrt{n}\rfloor$ centers. If we get too many centers, we will reset the net with a larger radius. Each block may be assigned to a center, such that every center is assigned at most one block. Initially, every block is unassigned. Generally, if a newly received point lies within the ball of a center, it will be placed inside the block assigned to that center. The exact algorithm is given below as Fill-Most-Blocks.
@@
Fill-Most-Blocks($n,A,X$) 1. Let $N_{1}=\lfloor\sqrt{n}\rfloor$ and $N_{2}=2N_{1}$. 2. Partition $A$ into $N_{2}$ subarrays, called blocks, of length at least $\lfloor n/N_{2}\rfloor$. 3. Initialize $r=0$ and $C=\emptyset$. 4. For each point $x$ in the stream $X$: (a) Increase-Net$(C,r,x)$. (b) If $|C|>N_{1}$: i. Unassign all blocks. ii. Set $r=4\operatorname{\mathtt{MST}}(X^{\prime})/N_{1}$, where $X^{\prime}$ is the set of known input points. iii. Set $C=\{x\}$. (c) Let $c\in C$ be a center with $d(x,c)\leq r$. (d) If a full block $B$ is assigned to $c$: i. Unassign $B$ from $c$. (e) If no block is assigned to $c$: i. Assign an unassigned non-full block to $c$. (f) Let $B$ be the block assigned to $c$. (g) Place $x$ in the left-most empty cell of $B$.
@@
Claim. After step 4b, $C$ is an $r$-net of a set containing $x$.
Proof of claim. Clearly, the claim holds if step 4(b)iii just ran, so let us show, that it also holds before and between these resets of the net. This follows by induction, where the base case is step 3 producing an $r$-net, and the induction step is step 4a with Lemma 8. $\triangleleft$
@@
Consider first the case where the net was not reset (steps 4(b)i to 4(b)iii) between $x$ and $x^{\prime}$ being placed. Then $B$ was assigned to the same center $c$ with the same radius $r$ both when $x$ and $x^{\prime}$ were placed. Let $X^{\prime}$ be the set of points read right after both $x$ and $x^{\prime}$ were placed. By the triangle inequality, $\operatorname{\mathtt{OPT}}(X^{\prime})\leq\operatorname{\mathtt{OPT}}(X)$, so by lemma 5, $d(x,x^{\prime})\leq d(x,c)+d(c,x^{\prime})\leq 2r=8\operatorname{\mathtt{MST}}(X^{\prime})/N_{1}\leq 8\operatorname{\mathtt{OPT}}(X^{\prime})/N_{1}\leq 8\operatorname{\mathtt{OPT}}(X)/N_{1}$. Since at most $\lceil n/2\rceil$ points are placed, the total cost between such pairs of points is at most $(\lceil n/2\rceil-1)8\operatorname{\mathtt{OPT}}(X)/N_{1}\leq 4n\operatorname{\mathtt{OPT}}(X)/N_{1}$.
Consider now the case where the net was reset between $x$ and $x^{\prime}$ being placed. Then $d(x,x^{\prime})\leq\operatorname{\mathtt{MST}}(X^{\prime})$ where $X^{\prime}$ is the set of points read at any point after placing both $x$ and $x^{\prime}$. This is a worse bound, but this case can only appear once per block per reset of the net. So the total cost for such pairs of points is at most $N_{2}\operatorname{\mathtt{MST}}(X^{\prime})$ for each reset, where $X^{\prime}$ is the set of points read when resetting. Let $X_{1}$ be the set of points read at the point of a reset, or at the beginning of the algorithm, and $X_{2}$ be the set of points read at the point of the following reset or when the algorithm terminates. Right before the latter reset, $|C|>N_{1}$ and $r=4\operatorname{\mathtt{MST}}(X_{1})/N_{1}$. This is true even if $X_{1}$ is the set at the beginning of the algorithm, as then $X_{1}=\emptyset$ and $r=0$. So Lemma 7 gives us that $2\operatorname{\mathtt{MST}}(X_{2})\geq(|C|-1)r\geq N_{1}4\operatorname{\mathtt{MST}}(X_{1})/N_{1}=4\operatorname{\mathtt{MST}}(X_{1})$, which simplifies to $\operatorname{\mathtt{MST}}(X_{2})\geq 2\operatorname{\mathtt{MST}}(X_{1})$. The total cost of this type across all rebuilds is thus no more than the geometric series $N_{2}\sum_{i=0}2^{-i}\operatorname{\mathtt{MST}}(X)=2N_{2}\operatorname{\mathtt{MST}}(X)\leq 2N_{2}\operatorname{\mathtt{OPT}}(X)$, where the last inequality uses Lemma 5.
@@ 4.1 Impossibility of optimality for every fixed metric space
By closely following the proof of Theorem 11, replacing their random input by $\mathcal{X}$, we get that the expected cost of any deterministic algorithm on $\mathcal{X}$ is $\Omega(n)$. We provide a full proof in Section A.2, restating this claim as Lemma 13.
@@
By assumption, $\mathcal{A}$ has competitive ratio $O(\sqrt{n})$ in general, and competitive ratio $O(\log n)$ when $x$ is not in the input. It’s easy to see that $\operatorname{\mathtt{OPT}}(\mathcal{X})\in O(n^{4/5})$, so we have $\operatorname{\mathbb{E}}[\mathcal{A}(\mathcal{X})\mid x\in\mathcal{X}]\in O(n^{4/5}\sqrt{n})$ and $\operatorname{\mathbb{E}}[\mathcal{A}(\mathcal{X})\mid x\not\in\mathcal{X}]\in O(n^{4/5}\log n)$. In total, $\operatorname{\mathbb{E}}[\mathcal{A}(\mathcal{X})]\in O(n^{-2/5}n^{4/5}\sqrt{n}+n^{4/5}\log n)\subseteq o(n)$, contradicting $\operatorname{\mathbb{E}}[\mathcal{A}(\mathcal{X})]\in\Omega(n)$. ∎